OpenStudy (anonymous):

write a parametric equation for a circle with a radius of 20 cm and when t=0sec the x,y coordinates are (20,0) the circle takes 3 sec to complete one revolution

OpenStudy (dumbcow):

period = 3 B = 2pi/3 \[x(t) = 20\cos (\frac{2\pi}{3}t)\] \[y(t) = 20\sin (\frac{2\pi}{3}t)\]